# CryoGS: Gaussian Splatting for Cryo-EM Homogeneous Reconstruction

### v1

@@ Abstract
As a critical modality for structural biology, cryogenic electron microscopy (cryo-EM) facilitates the determination of macromolecular structures at near-atomic resolution. The core computational task in single-particle cryo-EM is to reconstruct the 3D electrostatic potential of a molecule from a large collection of noisy 2D projections acquired at unknown orientations. Gaussian mixture models (GMMs) provide a continuous, compact, and physically interpretable representation for molecular density and have recently gained interest in cryo-EM reconstruction. However, existing methods rely on external consensus maps or atomic models for initialization, limiting their use in self-contained pipelines. Addressing this issue, we introduce cryoGS, a GMM-based method that integrates Gaussian splatting with the physics of cryo-EM image formation. In particular, we develop an orthogonal projection-aware Gaussian splatting, with adaptations such as a normalization term and FFT-aligned coordinate system tailored for cryo-EM imaging. All these innovations enable stable and efficient homogeneous reconstruction directly from raw cryo-EM particle images using random initialization. Experimental results on real datasets validate the effectiveness and robustness of cryoGS over representative baselines. The code will be released upon publication.
@@ Introduction
In this work, we propose cryoGS, a GMM-based homogeneous reconstruction method that fills this foundational gap. Given known particle poses, cryoGS performs stable and efficient reconstruction directly from raw cryo-EM projection images, starting from random initialization and requiring no external priors. Inspired by recent advances in 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023), we model the 3D density as a mixture of anisotropic Gaussians and project them into 2D using a novel differentiable orthographic splatting algorithm consistent with cryo-EM physics. To support practical scalability and training efficiency, we develop a CUDA-accelerated real-space renderer that enables fast rasterization and optimization of the GMM.
@@
  • A self-contained GMM-based reconstruction method: We present cryoGS as the first method capable of performing cryo-EM homogeneous reconstruction from scratch using only a Gaussian Mixture Model, thereby establishing the missing foundation needed to develop GMMs into standalone reconstruction tools.
@@
  • Experimental validation: We demonstrate the effectiveness of cryoGS on real datasets, showing it converges reliably from random initialization and achieves reconstruction quality outperforming state-of-the-art methods.
@@
Figure 1: Cryo-EM reconstruction aims to recover a 3D volume (a) from a large set of 2D particle images (b). (b) Purified biomolecules with random orientations are embedded in a thin layer of vitreous ice. The electrostatic potential of the sample interacts with transmitted electrons, forming a micrograph that contains 2D projections of the molecules. Individual particle images are extracted from the micrograph; they are extremely noisy and modulated by highly oscillatory CTFs. (c) Fourier slice theorem: the 2D Fourier transform of a particle image corresponds to a central slice of the 3D Fourier transform of the volume. (d) Overview of cryoGS. An anisotropic GMM representing the 3D volume is transformed to the projection direction, orthogonally projected onto a 2D image plane using a fast differentiable rasterizer, and modulated by the oscillatory CTF to simulate a physically accurate projection. The GMM parameters are optimized via gradients propagated from the discrepancy between the simulated and observed particle images. The resulting GMM can be voxelized to obtain the final 3D volume.
@@
Our goal is to achieve physically accurate and computationally efficient cryo-EM reconstruction by leveraging a Gaussian Mixture Model (GMM). To this end, we propose cryoGS, a differentiable framework that represents the 3D electrostatic potential of a specimen as a set of anisotropic Gaussians and directly simulates the cryo-EM image formation process in real space, faithfully adhering to the physics of transmission electron microscopy.
@@
To address these issues, cryoGS introduces several key adaptations: (i) we replace the heuristic alpha compositing with physically grounded line integrals to reflect the transmission nature of electron imaging; (ii) we fix the normalization between 3D-to-2D transformation and apply consistent learning rates across all parameters to ensure stable optimization; and (iii) we align the rasterization coordinate system with the FFT grid, allowing accurate gradient propagation through CTF modulation.
These modifications collectively enable cryoGS to perform stable, end-to-end differentiable reconstruction from raw cryo-EM particle images, starting from random initialization without relying on externally provided volumes.
@@
By substituting Eq. (1), we obtain the full forward process of cryoGS. Specifically, we apply a viewing transformation to align the GMM to the target orientation, orthographically project each Gaussian along the z-axis to form a 2D image, and convolve the result with the PSF:
@@
Figure 2: Qualitative and quantitative comparison of voxel-based, neural, and GMM-based representations. (Left) Final 3D reconstructions on four real datasets visualized with ChimeraX (Pettersen et al. 2021). (Right) FSC curves are plotted for quantitative evaluation. Gray dashed lines indicate the standard resolution thresholds of $0.5$ and $0.143$, reported in Angstroms (Å). CryoGS consistently achieves higher resolution across all datasets.
@@
We evaluate the performance of different volume representations on real cryo-EM datasets under a homogeneous reconstruction setting. To ensure a fair comparison focused solely on the choice of volume representation, all methods reconstruct consensus maps using the same set of published particle poses, without performing pose estimation. Our evaluation focuses on two aspects: (i) the ability to reconstruct high-resolution consensus maps, and (ii) robustness to noise and imperfect pose assignments. We compare voxel-based backprojection, the neural representation method cryoDRGN (Zhong et al. 2021a), and our proposed GMM-based method cryoGS. Visualizations of the reconstructed volumes are shown in Fig. 2, and spatial resolution is quantified using gold-standard FSC curves. Each volume of cryoGS is represented using 30,00030{,}000 Gaussians.
@@
The Pf80S ribosome is relatively easy to reconstruct due to its high-contrast images and structurally stable particles. All methods achieve high-resolution results ($3.80$ Å) and strong FSC agreement across the spectrum. cryoDRGN yields slightly higher FSC values at intermediate frequencies, while cryoGS outperforms all baselines at high spatial frequencies, demonstrating its ability to recover fine structural details.
@@
The RAG complex poses greater challenges due to imperfect published poses and flexible regions such as the DNA elements and the nonamer binding domain (NBD), indicated by arrows. CryoGS significantly outperforms both backprojection and cryoDRGN, achieving a resolution of $2.49$ Å. Unlike the baselines, cryoGS reconstructs the DNA elements and the NBD with minimal artifacts. Its FSC curve remains consistently above those of other methods across all spatial frequencies, highlighting its robustness to pose inaccuracies and structural variability.
@@
This LSU assembly dataset contains substantial compositional and conformational heterogeneity, making consensus reconstruction particularly challenging. Both FSC analysis and visualization show that cryoGS is more resilient under such conditions, achieving a resolution of $3.30$ Å with fewer artifacts than voxel-based or neural methods, as indicated by the red circle.
@@
The spliceosome dataset features large-scale motions of the SF3b indicated by the red circle, making consensus reconstruction particularly challenging. The reconstructions from backprojection and cryoDRGN show pronounced artifacts in this region, while cryoGS is more robust to such motions and achieves a resolution of $4.26$ Å. FSC analysis further confirms that cryoGS significantly outperforms the baselines across the frequency range.
CryoGS consistently converges within just $5$ epochs, with FSC curves from the $4$th and $5$th epochs tightly overlapping, indicating stable optimization and improved generalization in consensus reconstruction.
@@
Fig. 3 shows the FSC curves for cryoGS with varying numbers of Gaussians. In general, increasing the number of Gaussians leads to improved FSC, as a denser GMM provides greater representational capacity. While cryoGS performs well on most datasets, its relative performance varies due to differences in structural complexity and dataset-specific challenges. On EMPIAR-10028, cryoGS reaches a resolution of $3.8$ Å under all settings. While configurations with fewer than 10,00010{,}000 Gaussians exhibit lower FSC values than cryoDRGN and backprojection across most frequencies, the curves intersect at the highest frequency, indicating comparable final resolution. For EMPIAR-10049, all cryoGS settings significantly outperform both backprojection ($4.00$ Å) and cryoDRGN ($4.07$ Å), achieving a resolution of $2.49$ Å. Moreover, the FSC curves of all cryoGS variants remain consistently above those of the two baselines across the entire frequency range. For EMPIAR-10076, the 30,00030{,}000-Gaussian model clearly outperforms other settings; even with fewer Gaussians, cryoGS still surpasses the baselines, reaching $3.3$ Å. For EMPIAR-10180, the models with 10,00010{,}000 and 30,00030{,}000 Gaussians achieve the best FSC, reaching $4.3$ Å, while sparser GMMs remain competitive at high spatial frequencies. Overall, we find that using 10,00010{,}000 Gaussians is sufficient to reconstruct high-resolution volumes that consistently outperform baseline methods across most datasets.
@@ Conclusion
We present cryoGS, a novel GMM-based framework that integrates Gaussian splatting with the physics of cryo-EM image formation. CryoGS enables stable and efficient homogeneous reconstruction directly from raw cryo-EM particle images, starting from random initialization without relying on consensus volumes. Experimental results on real datasets demonstrate the effectiveness, robustness, and faster convergence of cryoGS compared to representative baselines.
Limitation and future work. While our current method assumes known poses and thus does not qualify as an ab initio approach, cryoGS establishes a principled foundation for future GMM-based methods that aim to tackle ab initio and heterogeneous reconstruction. We believe cryoGS provides a missing piece in the broader goal of integrating GMMs into scalable and self-contained cryo-EM reconstruction pipelines. These directions are left for future work.
@@ Appendix C More Implementation Details
CryoGS (Ours) $11\cdot N$ N=2,048N=2{,}048 11 $344$ $346$
@@
Table S1: Comparison of GPU memory usage across reconstruction methods at $D=192$ and $D=256$. CryoGS achieves the lowest and most stable memory footprint, while cryoDRGN incurs significantly higher usage due to its deep neural architecture. All measurements were performed on a single NVIDIA RTX 3090, reflecting the memory usage during the reconstruction of a single full density map. Memory for voxelization and FSC-related half-map reconstruction is not included.
@@ Appendix D Additional Results
Tab. S1 presents a comparison of GPU memory usage across different reconstruction methods. CryoDRGN (Zhong et al. 2021a) exhibits the highest memory footprint, exceeding $2.5$ GiB at $D=192$ and approaching $5$ GiB at $D=256$, primarily due to its deep neural decoder and a larger batch size of 8. Interestingly, backprojection consumes more memory at $D=192$ than at $D=256$, which may be attributed to implementation-specific factors such as padding overhead or kernel-level optimizations that favor power-of-two dimensions. This anomaly appears method-specific and does not reflect a general trend. In contrast, cryoGS demonstrates consistently low memory usage across all configurations. Even with as many as 30,00030{,}000 Gaussians, cryoGS maintains a GPU memory footprint below $380$ MiB, with negligible variation across resolutions. This efficiency underscores the scalability and suitability of cryoGS for large-scale or memory-constrained cryo-EM reconstruction scenarios.
@@
We investigate how the choice between isotropic and anisotropic Gaussians affects reconstruction quality. While cryoGS adopts anisotropic Gaussians to represent 3D volumes, it remains compatible with the isotropic formulation commonly used in prior work (2021; 2023; 2023; 2024; 2025). As shown in Fig. S1, anisotropic GMMs achieve higher FSC scores across spatial frequencies and produce sharper, more detailed structures. Subjectively, isotropic Gaussians struggle to capture directionally extended structures and are more prone to noise, often resulting in oversmoothing or distortion. This geometric rigidity may impair their ability to fit complex structures during early optimization, contributing to the unstable convergence observed with random initialization in prior methods. These results highlight the improved representational capacity and reconstruction robustness enabled by anisotropic modeling.

### v2

# CryoSplat: Gaussian Splatting for Cryo-EM Homogeneous Reconstruction

@@ Abstract
As a critical modality for structural biology, cryogenic electron microscopy (cryo-EM) facilitates the determination of macromolecular structures at near-atomic resolution. The core computational task in single-particle cryo-EM is to reconstruct the 3D electrostatic potential of a molecule from a large collection of noisy 2D projections acquired at unknown orientations. Gaussian mixture models (GMMs) provide a continuous, compact, and physically interpretable representation for molecular density and have recently gained interest in cryo-EM reconstruction. However, existing methods rely on external consensus maps or atomic models for initialization, limiting their use in self-contained pipelines. Addressing this issue, we introduce cryoSplat, a GMM-based method that integrates Gaussian splatting with the physics of cryo-EM image formation. In particular, we develop an orthogonal projection-aware Gaussian splatting, with adaptations such as a normalization term and FFT-aligned coordinate system tailored for cryo-EM imaging. All these innovations enable stable and efficient homogeneous reconstruction directly from raw cryo-EM particle images using random initialization. Experimental results on real datasets validate the effectiveness and robustness of cryoSplat over representative baselines. The code will be released upon publication.
@@ Introduction
In this work, we propose cryoSplat, a GMM-based homogeneous reconstruction method that fills this foundational gap. Given known particle poses, cryoSplat performs stable and efficient reconstruction directly from raw cryo-EM projection images, starting from random initialization and requiring no external priors. Inspired by recent advances in 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023), we model the 3D density as a mixture of anisotropic Gaussians and project them into 2D using a novel differentiable orthographic splatting algorithm consistent with cryo-EM physics. To support practical scalability and training efficiency, we develop a CUDA-accelerated real-space renderer that enables fast rasterization and optimization of the GMM.
@@
  • A self-contained GMM-based reconstruction method: We present cryoSplat as the first method capable of performing cryo-EM homogeneous reconstruction from scratch using only a Gaussian Mixture Model, thereby establishing the missing foundation needed to develop GMMs into standalone reconstruction tools.
@@
  • Experimental validation: We demonstrate the effectiveness of cryoSplat on real datasets, showing it converges reliably from random initialization and achieves reconstruction quality outperforming state-of-the-art methods.
@@
Figure 1: Cryo-EM reconstruction aims to recover a 3D volume (a) from a large set of 2D particle images (b). (b) Purified biomolecules with random orientations are embedded in a thin layer of vitreous ice. The electrostatic potential of the sample interacts with transmitted electrons, forming a micrograph that contains 2D projections of the molecules. Individual particle images are extracted from the micrograph; they are extremely noisy and modulated by highly oscillatory CTFs. (c) Fourier slice theorem: the 2D Fourier transform of a particle image corresponds to a central slice of the 3D Fourier transform of the volume. (d) Overview of cryoSplat. An anisotropic GMM representing the 3D volume is transformed to the projection direction, orthogonally projected onto a 2D image plane using a fast differentiable rasterizer, and modulated by the oscillatory CTF to simulate a physically accurate projection. The GMM parameters are optimized via gradients propagated from the discrepancy between the simulated and observed particle images. The resulting GMM can be voxelized to obtain the final 3D volume.
@@
Our goal is to achieve physically accurate and computationally efficient cryo-EM reconstruction by leveraging a Gaussian Mixture Model (GMM). To this end, we propose cryoSplat, a differentiable framework that represents the 3D electrostatic potential of a specimen as a set of anisotropic Gaussians and directly simulates the cryo-EM image formation process in real space, faithfully adhering to the physics of transmission electron microscopy.
@@
To address these issues, cryoSplat introduces several key adaptations: (i) we replace the heuristic alpha compositing with physically grounded line integrals to reflect the transmission nature of electron imaging; (ii) we fix the normalization between 3D-to-2D transformation and apply consistent learning rates across all parameters to ensure stable optimization; and (iii) we align the rasterization coordinate system with the FFT grid, allowing accurate gradient propagation through CTF modulation.
These modifications collectively enable cryoSplat to perform stable, end-to-end differentiable reconstruction from raw cryo-EM particle images, starting from random initialization without relying on externally provided volumes.
@@
By substituting Eq. (1), we obtain the full forward process of cryoSplat. Specifically, we apply a viewing transformation to align the GMM to the target orientation, orthographically project each Gaussian along the z-axis to form a 2D image, and convolve the result with the PSF:
@@
Figure 2: Qualitative and quantitative comparison of voxel-based, neural, and GMM-based representations. (Left) Final 3D reconstructions on four real datasets visualized with ChimeraX (Pettersen et al. 2021). (Right) FSC curves are plotted for quantitative evaluation. Gray dashed lines indicate the standard resolution thresholds of $0.5$ and $0.143$, reported in Angstroms (Å). CryoSplat consistently achieves higher resolution across all datasets.
@@
We evaluate the performance of different volume representations on real cryo-EM datasets under a homogeneous reconstruction setting. To ensure a fair comparison focused solely on the choice of volume representation, all methods reconstruct consensus maps using the same set of published particle poses, without performing pose estimation. Our evaluation focuses on two aspects: (i) the ability to reconstruct high-resolution consensus maps, and (ii) robustness to noise and imperfect pose assignments. We compare voxel-based backprojection, the neural representation method cryoDRGN (Zhong et al. 2021a), and our proposed GMM-based method cryoSplat. Visualizations of the reconstructed volumes are shown in Fig. 2, and spatial resolution is quantified using gold-standard FSC curves. Each volume of cryoSplat is represented using 30,00030{,}000 Gaussians.
@@
The Pf80S ribosome is relatively easy to reconstruct due to its high-contrast images and structurally stable particles. All methods achieve high-resolution results ($3.80$ Å) and strong FSC agreement across the spectrum. cryoDRGN yields slightly higher FSC values at intermediate frequencies, while cryoSplat outperforms all baselines at high spatial frequencies, demonstrating its ability to recover fine structural details.
@@
The RAG complex poses greater challenges due to imperfect published poses and flexible regions such as the DNA elements and the nonamer binding domain (NBD), indicated by arrows. CryoSplat significantly outperforms both backprojection and cryoDRGN, achieving a resolution of $2.49$ Å. Unlike the baselines, cryoSplat reconstructs the DNA elements and the NBD with minimal artifacts. Its FSC curve remains consistently above those of other methods across all spatial frequencies, highlighting its robustness to pose inaccuracies and structural variability.
@@
This LSU assembly dataset contains substantial compositional and conformational heterogeneity, making consensus reconstruction particularly challenging. Both FSC analysis and visualization show that cryoSplat is more resilient under such conditions, achieving a resolution of $3.30$ Å with fewer artifacts than voxel-based or neural methods, as indicated by the red circle.
@@
The spliceosome dataset features large-scale motions of the SF3b indicated by the red circle, making consensus reconstruction particularly challenging. The reconstructions from backprojection and cryoDRGN show pronounced artifacts in this region, while cryoSplat is more robust to such motions and achieves a resolution of $4.26$ Å. FSC analysis further confirms that cryoSplat significantly outperforms the baselines across the frequency range.
CryoSplat consistently converges within just $5$ epochs, with FSC curves from the $4$th and $5$th epochs tightly overlapping, indicating stable optimization and improved generalization in consensus reconstruction.
@@
Fig. 3 shows the FSC curves for cryoSplat with varying numbers of Gaussians. In general, increasing the number of Gaussians leads to improved FSC, as a denser GMM provides greater representational capacity. While cryoSplat performs well on most datasets, its relative performance varies due to differences in structural complexity and dataset-specific challenges. On EMPIAR-10028, cryoSplat reaches a resolution of $3.8$ Å under all settings. While configurations with fewer than 10,00010{,}000 Gaussians exhibit lower FSC values than cryoDRGN and backprojection across most frequencies, the curves intersect at the highest frequency, indicating comparable final resolution. For EMPIAR-10049, all cryoSplat settings significantly outperform both backprojection ($4.00$ Å) and cryoDRGN ($4.07$ Å), achieving a resolution of $2.49$ Å. Moreover, the FSC curves of all cryoSplat variants remain consistently above those of the two baselines across the entire frequency range. For EMPIAR-10076, the 30,00030{,}000-Gaussian model clearly outperforms other settings; even with fewer Gaussians, cryoSplat still surpasses the baselines, reaching $3.3$ Å. For EMPIAR-10180, the models with 10,00010{,}000 and 30,00030{,}000 Gaussians achieve the best FSC, reaching $4.3$ Å, while sparser GMMs remain competitive at high spatial frequencies. Overall, we find that using 10,00010{,}000 Gaussians is sufficient to reconstruct high-resolution volumes that consistently outperform baseline methods across most datasets.
@@ Conclusion
We present cryoSplat, a novel GMM-based framework that integrates Gaussian splatting with the physics of cryo-EM image formation. CryoSplat enables stable and efficient homogeneous reconstruction directly from raw cryo-EM particle images, starting from random initialization without relying on consensus volumes. Experimental results on real datasets demonstrate the effectiveness, robustness, and faster convergence of cryoSplat compared to representative baselines.
Limitation and future work. While our current method assumes known poses and thus does not qualify as an ab initio approach, cryoSplat establishes a principled foundation for future GMM-based methods that aim to tackle ab initio and heterogeneous reconstruction. We believe cryoSplat provides a missing piece in the broader goal of integrating GMMs into scalable and self-contained cryo-EM reconstruction pipelines. These directions are left for future work.
@@ Appendix C More Implementation Details
CryoSplat (Ours) $11\cdot N$ N=2,048N=2{,}048 11 $344$ $346$
@@
Table S1: Comparison of GPU memory usage across reconstruction methods at $D=192$ and $D=256$. CryoSplat achieves the lowest and most stable memory footprint, while cryoDRGN incurs significantly higher usage due to its deep neural architecture. All measurements were performed on a single NVIDIA RTX 3090, reflecting the memory usage during the reconstruction of a single full density map. Memory for voxelization and FSC-related half-map reconstruction is not included.
@@ Appendix D Additional Results
Tab. S1 presents a comparison of GPU memory usage across different reconstruction methods. CryoDRGN (Zhong et al. 2021a) exhibits the highest memory footprint, exceeding $2.5$ GiB at $D=192$ and approaching $5$ GiB at $D=256$, primarily due to its deep neural decoder and a larger batch size of 8. Interestingly, backprojection consumes more memory at $D=192$ than at $D=256$, which may be attributed to implementation-specific factors such as padding overhead or kernel-level optimizations that favor power-of-two dimensions. This anomaly appears method-specific and does not reflect a general trend. In contrast, cryoSplat demonstrates consistently low memory usage across all configurations. Even with as many as 30,00030{,}000 Gaussians, cryoSplat maintains a GPU memory footprint below $380$ MiB, with negligible variation across resolutions. This efficiency underscores the scalability and suitability of cryoSplat for large-scale or memory-constrained cryo-EM reconstruction scenarios.
@@
We investigate how the choice between isotropic and anisotropic Gaussians affects reconstruction quality. While cryoSplat adopts anisotropic Gaussians to represent 3D volumes, it remains compatible with the isotropic formulation commonly used in prior work (2021; 2023; 2023; 2024; 2025). As shown in Fig. S1, anisotropic GMMs achieve higher FSC scores across spatial frequencies and produce sharper, more detailed structures. Subjectively, isotropic Gaussians struggle to capture directionally extended structures and are more prone to noise, often resulting in oversmoothing or distortion. This geometric rigidity may impair their ability to fit complex structures during early optimization, contributing to the unstable convergence observed with random initialization in prior methods. These results highlight the improved representational capacity and reconstruction robustness enabled by anisotropic modeling.